A bicycle wheel with a radius of 0.40 meters has a
Answer #1

24) d) -3.06 rad/s^2

alfa = (wf-wi)/t

= (0 - 15.3)/5

= -3.06 rad/s^2

25) e) 0.36 kg.m^2

I = M*R^2

= 2.25*0.4^2

= 0.36 kg.m^2

26) d)42.13 J

KEi = 0.5*I*wi^2

= 0.5*0.36*15.3^2

= 42.13 J

27) a) 1.01 N.m

Apply, |T| = I*alfa

==> T = 0.36*3.06

= 1.1016 N.m